OpenStudy (anonymous):

Find y' given tan-1(xy) = 1 + x2y

OpenStudy (anonymous):

x2y= x^2y

OpenStudy (across):

This is a good example of implicit differentiation!

OpenStudy (anonymous):

implicit diff for this one \[\frac{d}{dx}\tan^{-1}(xy)=\frac{1}{(xy)^2+1}\times y'\] by the chain rule

OpenStudy (anonymous):

similarly \[\frac{d}{dx}[1+x^2y]=2xy+x^2y'\] again by product and chain rule. set them equal and solve for \[y'\] algebra from here on in

OpenStudy (mr.math):

I hope I didn't make any mistakes in my simplification.

OpenStudy (anonymous):

bty if that is not clear, think \[y=f(x),y'=f'(x)\] and imagine how you would take the derivative of \[1+x^2\times f(x)\] and also \[\tan^{-1}(xf(x))\]

OpenStudy (anonymous):

the algebra is the annoying part. i like to cheat http://www.wolframalpha.com/input/?i=arctan%28xy%29+%3D+1%2Bx^2y

OpenStudy (mr.math):

I did make a mistake :(

OpenStudy (anonymous):

i didn't, because i didn't do it

OpenStudy (mr.math):

My last step is not correct @indianalexander.

OpenStudy (mr.math):

@satellite: It's hard to do these operation on the keyboard, I need a pen and a paper :)

OpenStudy (mr.math):

Wait! wait! \[\frac{d}{dx}\tan^{-1}(xy)=\frac{xy'+y}{1+x^2y^2}.\]

OpenStudy (mr.math):

We made a mistake above. You should set that equal to \(x^2y'+2xy\).

OpenStudy (mr.math):

Then simplify!